OpenStudy (anonymous):

Please help! Prove: cos θ - cos θ•sin2 θ = cos3 θ. You must show all work.

OpenStudy (anonymous):

which are exponents and which are not? ARe there any exponents ?

OpenStudy (anonymous):

like sin^2x or sin(2x) ?

OpenStudy (anonymous):

im not sure :/

OpenStudy (anonymous):

\(\Large\color{black}{ \bf cos θ - cos θ~sin^2 θ = cos^3 θ }\) like this ?

OpenStudy (anonymous):

there is supposed to be a multiplication in between cos0 and sin^20

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

When there is no space that's what it means (just like "ab" mean a times b ) Anyway, .... \(\Large\color{black}{ \bf cos θ - cosθ~sin^2 θ= cos^3 θ }\) factor out of cos θ on the left \(\Large\color{black}{ \bf cos θ(1 - ~sin^2 θ) = cos^3 θ }\)

OpenStudy (anonymous):

\(\Large\color{black}{ \bf cos θ(1 - ~sin^2 θ) = cos^3 θ }\) Divide both sides by cos θ, \(\Large\color{black}{ \bf 1 - ~sin^2 θ = cos^2 θ }\) add sin^2 θ to both sides \(\Large\color{black}{ \bf 1= cos^2 θ+ ~sin^2 θ }\)

OpenStudy (anonymous):

thankyou so much, would that be my final answer or do i still need to add?

OpenStudy (anonymous):

No, that is considered to be authentic, although, there is a proof to it as well.

OpenStudy (anonymous):

thankyou!!